OpenStudy (anonymous):

Solve. 2x2 + 13x = –20

mathslover (mathslover):

ok lemme give you a brief intro. to quadratic equation : \[\large{ax^2+bx+c=0}\] -> general form of a quadratic equation

mathslover (mathslover):

what do you think is the above equation , a quadratic equation?

OpenStudy (anonymous):

Yes

mathslover (mathslover):

ok good , so remember that the roots of a quadratic equation are : \[\large{roots =\frac{-b\pm \sqrt{b^2-4ac}}{2a}}\]

mathslover (mathslover):

put b = 13 , a = 2 and c = 20

OpenStudy (anonymous):

I'm not sure

mathslover (mathslover):

ok .. see solve this now : \[\large{\frac{-13\pm \sqrt{169-160}}{4}}\] \[\large{\frac{-13\pm \sqrt{9}}{4}}\]

OpenStudy (anonymous):

x= 1/3 and x=3?

mathslover (mathslover):

-10/4 = -5/2 and -16/4 = -4

mathslover (mathslover):

sqrt{9}=3

OpenStudy (anonymous):

Ohhh. Wait okay I see what I did wrong. Thank you!